# Harnessing the Power of Neural Operators with Automatically Encoded Conservation Laws

### v2

@@ Abstract
Neural operators (NOs) have emerged as effective tools for modeling complex physical systems in scientific machine learning. In NOs, a central characteristic is to learn the governing physical laws directly from data. In contrast to other machine learning applications, partial knowledge is often known a priori about the physical system at hand whereby quantities such as mass, energy and momentum are exactly conserved. Currently, NOs have to learn these conservation laws from data and can only approximately satisfy them due to finite training data and random noise. In this work, we introduce conservation law-encoded neural operators (clawNOs), a suite of NOs that endow inference with automatic satisfaction of such conservation laws. ClawNOs are built with a divergence-free prediction of the solution field, with which the continuity equation is automatically guaranteed. As a consequence, clawNOs are compliant with the most fundamental and ubiquitous conservation laws essential for correct physical consistency. As demonstrations, we consider a wide variety of scientific applications ranging from constitutive modeling of material deformation, incompressible fluid dynamics, to atmospheric simulation. ClawNOs significantly outperform the state-of-the-art NOs in learning efficacy, especially in small-data regimes.
@@ 1 Introduction
To improve learning efficacy and robustness in small-data regimes, we propose to encode a series of fundamental conservation laws into the architecture of NOs. Their inference is then constrained to a physically-consistent manifold. Here, we focus on the conservation of mass or volume which leads to a continuity equation for divergence-free flow. The development is based on two key innovations. Firstly, once the output function of a NO is divergence-free, the continuity equation is automatically satisfied, guaranteeing conservation of volume and mass. Based on the concept of differential forms, the conservation law is embedded through building divergence-free output functions. Secondly, to evaluate the differential forms, we propose an additional linear layer in NOs, whose weights are pre-specified based on high-order numerical differentiations on the given grids. Given an input function and its values on given grids, this layer evaluates the spatial derivatives as weighted linear combinations of neighboring or global points, mapping the function to its approximated derivatives.
Figure 1: Predictability demo in atmospheric modeling. While FNO plausibly learns the wave propagation patterns, its relative L2 error is almost twice as big as that of clawNO (cf. Section 4.3), especially in the vicinity of mountains (Fig. 5). ClawNOs automatically satisfy the conservation law and improve physical consistency.
Compared to existing NO methods, the proposed architecture mainly carries three significant advantages. First, different from existing physics-informed neural operators (Goswami et al., 2022a; b; Li et al., 2021; Wang et al., 2021), our approach is readily applicable to learn physical systems directly from experimental measurements, since it only requires observed data pairs and does not rely on fully known governing equations. Second, the conservation laws in our approach are realized through built-in architectures, which constrain the output function to a physically consistent manifold independent of noisy measurement and/or scarce data coverage. Third, our architecture is designed to encode general conservation laws at large. Besides the conservation of mass, linear/angular momentum as demonstrated in our examples, it is readily applicable to the conservations of energy and electric charge, and can be easily scaled up to higher dimensions. In summary, the main contributions of our work are:
@@
  • ClawNO only requires data pairs and does not rely on a priori domain knowledge, while the guaranteed conservation laws improve the learning efficacy, especially in small-data regime.
@@ 2 Background and related work
Learning hidden physics. Learning and predicting complex physics directly from data is ubiquitous in many scientific and engineering applications (Ghaboussi et al., 1998; 1991; Carleo et al., 2019; Karniadakis et al., 2021; Zhang et al., 2018; Cai et al., 2022; Pfau et al., 2020; He et al., 2021; Besnard et al., 2006). Amongst these real-world physical problems, the underlying governing laws largely remain unknown, and machine learning models act to discover hidden physics from data through training. Successful examples include graph neural networks (GNNs) in discovering molecular properties (Wieder et al., 2020; Wu et al., 2023a), neural ODEs in constructing strain energy formulation for materials (Tac et al., 2022), and CNNs in steady flow approximation (Guo et al., 2016), etc. In this work, we focus on learning hidden physics from measurements with fundamental physical laws enforced. Examples include learning material deformation model from experimental measurements, where the mass should be conserved, but the constitutive law remains hidden.
Neural operator learning. Among others, NOs possess superiority in discovering physical laws as function mappings. Contrary to classical neural networks (NNs) that operate between finite-dimensional Euclidean spaces, NOs are designed to learn mappings between infinite-dimensional function spaces (Li et al., 2020a; b; c; You et al., 2022a; Ong et al., 2022; Cao, 2021; Lu et al., 2019; 2021; Goswami et al., 2022a; Gupta et al., 2021). As a result, NOs are often employed to manifest the mapping between spatial and/or spatio-temporal data pairs. Compared with classical NNs, the most notable advantages of NOs are their resolution independence and generalizability to different input instances. Moreover, NOs only require data with no knowledge on the underlying governing laws. All these advantages make NOs promising tools to learn hidden physics directly from data.
Despite the aforementioned advances, purely data-driven NOs still suffer from the data challenge: they require a large set of paired data, which is prohibitively expensive in many engineering applications. To resolve this challenge, physics-informed neural operator (PINO) (Li et al., 2021) and physics-informed DeepONets (Goswami et al., 2022a; Wang et al., 2021) are introduced, where a PDE-based loss is added to the training loss as a penalization term. However, these approaches often require a priori knowledge of all the underlying physics laws (in the form of governing PDEs). Moreover, the penalization term only imposes the these knowledge as a weak constraint.
@@ 3 ClawNO: Conservation Law Encoded Neural Operator
We consider the learning of complex physical systems based on a number of observations of the input function values $\mathbf{f}_{i}(\mathbf{x})\in\mathbb{F}({\Omega};\mathbb{R}^{p_{f}})$ and the corresponding output function values $\mathbf{u}_{i}(\mathbf{x})\in\mathbb{U}({\Omega};\mathbb{R}^{p_{u}})$. Here, $i$ denotes the sample index, ${\Omega}\in\mathbb{R}^{p}$ is the bounded domain of interest, and $\mathbb{F}$ and $\mathbb{U}$ describe the Banach spaces of functions taking values in $\mathbb{R}^{p_{f}}$ and $\mathbb{R}^{p_{u}}$, respectively. To model the physical laws of such a system, we aim to learn the intrinsic operator $\mathcal{G}:\mathbb{F}\rightarrow\mathbb{U}$, that maps the input function $\mathbf{f}(\mathbf{x})$ to the output function $\mathbf{u}(\mathbf{x})$. In other words, given the measurements of $\mathbf{f}_{i}$ and $\mathbf{u}_{i}$ on a collection of points $\chi=\{\mathbf{x}_{j}\}^{M}_{j=1}\subset{\Omega}$, we seek to learn the physical response by constructing a parameterized surrogate operator of $\mathcal{G}$: $\tilde{\mathcal{G}}[\mathbf{f};\theta](\mathbf{x})\approx\mathbf{u}(\mathbf{x})$, where the trainable parameter set $\theta$ is obtained by solving the optimization problem:
$\min_{\theta\in\Theta}\mathcal{L}(\theta):=\min_{\theta\in\Theta}\sum_{i=1}^{N}\dfrac{1}{N}[C(\tilde{\mathcal{G}}[\mathbf{f}_{i};\theta],\mathbf{u}_{i})]\approx\min_{\theta\in\Theta}\dfrac{1}{N}\sum_{i=1}^{N}\dfrac{\sum_{j=1}^{M}{\left|\left|\tilde{\mathcal{G}}[\mathbf{f}_{i};\theta](\mathbf{x}_{j})-\mathbf{u}_{i}(\mathbf{x}_{j})\right|\right|}^{2}}{\sum_{j=1}^{M}{\left|\left|\mathbf{u}_{i}(\mathbf{x}_{j})\right|\right|}^{2}}.\vspace{-0.03in}$ (1)
@@
Although our approach is agnostic to any NO architecture, for the purpose of demonstration we consider the generic integral neural operators (Li et al., 2020b; c; You et al., 2022a; b) as the base models. In this context, an $L$-layer NO has the following form:
$\tilde{\mathcal{G}}[\mathbf{f};\theta](\mathbf{x}):=\mathcal{Q}\circ\mathcal{J}_{L}\circ\cdots\circ\mathcal{J}_{1}\circ\mathcal{P}[\mathbf{f}](\mathbf{x}),\vspace{-0.05in}$ (2)
@@
$\mathbf{h}(\mathbf{x},l+1)=\mathcal{J}^{FNO}_{l}[\mathbf{h}(\mathbf{x},l)]:=\sigma(W_{l}\mathbf{h}(\mathbf{x},l)+\mathcal{F}^{-1}[A_{l}\cdot\mathcal{F}[\mathbf{h}(\cdot,l)]](\mathbf{x})+\mathbf{c}_{l}),\vspace{-0.05in}$ (3)
@@ 3.1 Built-in divergence-free prediction
$\dfrac{\partial\rho}{\partial t}(\mathbf{x},t)+\nabla_{\mathbf{x}}\cdot\mu(\mathbf{x},t)=0,\quad\text{ for }(\mathbf{x},t)\in{\Omega}\times[0,T],\vspace{-0.05in}$ (5)
@@
where $\rho:\mathbb{R}^{p+1}\rightarrow\mathbb{R}^{+}$ is the volume density of the quantity to be conserved, $\mu:\mathbb{R}^{p+1}\rightarrow\mathbb{R}^{p}$ is the flux describing how this quantity flows, and $\nabla_{\mathbf{x}}$ denotes the gradient operator with respect to the spatial dimensions. Equation 5 states that the amount of the conserved quantity within a volume can only change by the amount of the quantity that flows in and out of the volume. Taking the mass conservation law as an instance, $\rho$ is the fluid’s density, and $\mathbf{u}$ is the velocity at each point, and then the mass flux becomes $\mu=\rho\mathbf{u}$. When considering the (quasi-)static problem, i.e., $\rho$ is independent of time, we have $\nabla_{\mathbf{x}}\cdot\mathbf{u}=0$, where the divergence is taken only over spatial variables. Such a divergence-free equation is also referred to as the condition of incompressibility in fluid dynamics. Moreover, we note that equation 5 can always be expressed as a divergence-free equation when considering both spatial and temporal dimensions. As a matter of fact, by taking $\mathbf{X}=(\mathbf{x},t)$ and $\mathbf{U}=(\mu,\rho)$, equation 5 can be re-written as $\nabla_{\mathbf{X}}\cdot\mathbf{U}=0$, there the gradient operator $\nabla_{\mathbf{X}}$ is taken with respect to both $\mathbf{x}$ and $t$. Then, the conservation law becomes equivalent to the static case, if we consider the augmented variable $\mathbf{X}$ on the augmented domain $\tilde{{\Omega}}:={\Omega}\times[0,T]$. Hence, in the following derivation we focus on the static case, with the goal of designing an NO architecture with the continuity equation $\text{div}(\mathbf{u}):=\nabla_{\mathbf{x}}(\mathbf{u})=0$ automatically guaranteed for its output function $\mathbf{u}$.
@@
To construct divergence-free NOs, we start from differential forms in $\mathbb{R}^{p}$, following the derivations in e.g., Barbarosie (2011); Kelliher (2021). More details are provided in Appendix A. Generally, an object that may be integrated over a $k-$dimensional manifold is called a $k-$form, which can be expressed as $\mu=\sum_{1\leq i_{1},i_{2},\cdots,i_{k}\leq p}\mu_{(i_{1},\cdots,i_{k})}dx_{i_{1}}\wedge\cdots\wedge dx_{i_{k}}:=\sum_{I}\mu_{I}dx_{I}$. For instance, any scalar function $g(\mathbf{x})\in C^{\infty}$ is a $0-$form, and $v=\sum_{i=1}^{p}v_{i}(\mathbf{x})dx_{i}$, $v_{i}(\mathbf{x})\in C^{\infty}$, is a $1-$form. Denote $d$ as the exterior derivative, which is an operation acting on a $k-$form and produces a $(k+1)-$form (e.g., $dg(\mathbf{x})=\sum_{i=1}^{p}\frac{\partial g}{\partial x_{i}}dx_{i}$), and the Hodge operator as $\star$, which matches to each $k-$form an $(p-k)-$form via $\star dx_{I}:=(-1)^{\sigma(I,I^{c})}dx_{I^{c}}$. Here, $I^{c}:=\{1,\cdots,n\}\backslash I$ is the complement set of $I$, and $\sigma(I,I^{c})$ is the sign of the permutation $(I,I^{c})$. A simple calculation yields that $\star\star\mu=(-1)^{k(p-k)}\mu$ and $dd\mu=0$ for any $k-$form $\mu$.
@@
To see the connection of these definitions to our goal, we note that our output function, $\mathbf{u}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$, can be equivalently expressed as a $1-$form $\mathbf{u}=\sum^{p}_{i=1}u_{i}(\mathbf{x})dx_{i}$. Then, the divergence of $\mathbf{u}$ can be expressed as $d\star\mathbf{u}$, and the above fundamental properties of the exterior derivative yields $\text{div}(\star d\mu)=d\star\star d\mu=dd\mu=0$ for any $(p-2)-$form $\mu$. That means, any vector field $\mathbf{u}$ satisfying $\mathbf{u}=\star d\mu$ is divergence-free. Our goal is therefore to parameterize the output function of NOs in the form of $\star d\mu$, or equivalently, as $\star d(\sum_{1\leq i_{1},\cdots,i_{p-2}\leq p}\mu_{(i_{1},\cdots,i_{p-2})}dx_{i_{1}}\wedge\cdots\wedge dx_{i_{p-2}}))=\sum_{i=1}^{p}\sum_{j=1}^{p}\frac{\partial\mu_{ij}}{\partial x_{j}}dx_{i}$. Here, $\mu_{ij}(\mathbf{x})$ is a function defined on ${\Omega}$ with $\mu_{ij}(\mathbf{x})=\mu_{(i,j)^{c}}$ for $i<j$ and $\mu_{ij}(\mathbf{x})=-\mu_{(j,i)^{c}}$ otherwise, as real-valued alternating functions. That means, taking any $\mu:=[\mu_{ij}(\mathbf{x})]$ satisfies $\mu_{ij}=-\mu_{ji}$ (a $p$ by $p$ skew-symmetric matrix-valued function), we have $\text{div}(\star d\mu)=0$, where $\star d\mu=[\text{div}(\mu_{1}),\cdots,\text{div}(\mu_{p})]^{T}$ and $\mu_{i}$ stands for the $i-$th row of $\mu$.

### v3

@@ Abstract
Neural operators (NOs) have emerged as effective tools for modeling complex physical systems in scientific machine learning. In NOs, a central characteristic is to learn the governing physical laws directly from data. In contrast to other machine learning applications, partial knowledge is often known a priori about the physical system at hand whereby quantities such as mass, energy and momentum are exactly conserved. Currently, NOs have to learn these conservation laws from data and can only approximately satisfy them due to finite training data and random noise. In this work, we introduce conservation law-encoded neural operators (clawNOs), a suite of NOs that endow inference with automatic satisfaction of such conservation laws. ClawNOs are built with a divergence-free prediction of the solution field, with which the continuity equation is automatically guaranteed. As a consequence, clawNOs are compliant with the most fundamental and ubiquitous conservation laws essential for correct physical consistency. As demonstrations, we consider a wide variety of scientific applications ranging from constitutive modeling of material deformation, incompressible fluid dynamics, to atmospheric simulation. ClawNOs significantly outperform the state-of-the-art NOs in learning efficacy, especially in small-data regimes. Our code and data accompanying this paper are available at https://github.com/ningliu-iga/clawNO.
Neural Operators, Conservation Laws, PDE Learning, Divergence-Free Prediction
@@ 1 Introduction
To improve learning efficacy and robustness in small-data regimes, we propose to encode a series of fundamental conservation laws into the architecture of NOs. Their inference is then constrained to a physically-consistent manifold. Here, we focus on the conservation of mass or volume which leads to a continuity equation for divergence-free flow. The development is based on two key innovations. Firstly, when the output function of a NO is divergence-free, the continuity equation is automatically satisfied, guaranteeing conservation of volume and mass. Based on the concept of differential forms, the conservation law is embedded through building divergence-free output functions. Secondly, to evaluate the differential forms, we propose an additional linear layer in NOs, whose weights are pre-specified based on high-order numerical differentiations on the given grids. Given an input function and its values on given grids, this layer evaluates the spatial derivatives as weighted linear combinations of neighboring or global points, mapping the function to its approximated derivatives.
Figure 1: Predictability demo in atmospheric modeling. While FNO plausibly learns the wave propagation patterns, its relative L2 error is twice as big as that of clawNO (cf. Section 4.3), especially in the vicinity of mountains (Fig. 6). ClawNOs automatically satisfy the conservation law and improve physical consistency.
Compared to existing NO methods, the proposed architecture mainly carries three significant advantages. First, different from existing physics-informed neural operators (Goswami et al., 2022a, b; Li et al., 2021; Wang et al., 2021), our approach is readily applicable to learn physical systems directly from experimental measurements, since it only requires observed data pairs and does not rely on fully known governing equations. Second, the conservation laws in our approach are realized through built-in architectures, which constrain the output function to a physically consistent manifold independent of noisy measurement and/or scarce data coverage. Third, our architecture is designed to encode general conservation laws at large. Besides the conservation of mass, linear/angular momentum as demonstrated in our examples, it is readily applicable to the conservations of energy and electric charge, and can be easily scaled up to higher dimensions. In summary, the main contributions of our work are:
@@
  • ClawNO only requires data pairs and does not rely on a priori domain knowledge, while the guaranteed conservation laws improve the learning efficacy, especially in the small-data regime.
@@ 2 Background and Related Work
Learning hidden physics. Learning and predicting complex physics directly from data is ubiquitous in many scientific and engineering applications (Ghaboussi et al., 1998, 1991; Carleo et al., 2019; Karniadakis et al., 2021; Zhang et al., 2018; Cai et al., 2022; Pfau et al., 2020; He et al., 2021; Besnard et al., 2006; Jafarzadeh et al., 2024; Liu et al., 2024b). Amongst these real-world physical problems, the underlying governing laws largely remain unknown, and machine learning models act to discover hidden physics from data through training. Successful examples include graph neural networks (GNNs) in discovering molecular properties (Wieder et al., 2020; Wu et al., 2023a), neural ODEs in constructing strain energy formulation for materials (Tac et al., 2022), and CNNs in steady flow approximation (Guo et al., 2016), etc. In this work, we focus on learning hidden physics from measurements with fundamental physical laws enforced. Examples include learning material deformation model from experimental measurements, where the mass should be conserved, but the constitutive law remains hidden.
Neural operator learning. Among others, NOs possess superiority in discovering physical laws as function mappings. Contrary to classical neural networks (NNs) that operate between finite-dimensional Euclidean spaces, NOs are designed to learn mappings between infinite-dimensional function spaces (Li et al., 2020a; Liu et al., 2024a; Li et al., 2020b, c; You et al., 2022a; Ong et al., 2022; Cao, 2021; Lu et al., 2019, 2021; Goswami et al., 2022a; Gupta et al., 2021). As a result, NOs are often employed to manifest the mapping between spatial and/or spatio-temporal data pairs. Compared with classical NNs, the most notable advantages of NOs are their resolution independence and generalizability to different input instances. Moreover, NOs only require data with no knowledge on the underlying governing laws. All these advantages make NOs promising tools to learn hidden physics directly from data.
Despite the aforementioned advances, purely data-driven NOs still suffer from the data challenge: they require a large set of paired data, which is prohibitively expensive in many engineering applications. To resolve this challenge, physics-informed neural operator (PINO) (Li et al., 2021) and physics-informed DeepONets (Goswami et al., 2022a; Wang et al., 2021) are introduced, where a PDE-based loss is added to the training loss as a penalization term. However, these approaches often require a priori knowledge of all the underlying physics laws (in the form of governing PDEs). Moreover, the penalization term only imposes these knowledge as a weak constraint.
@@ 3 ClawNO: Conservation Law Encoded Neural Operator
We consider the learning of complex physical systems based on a number of observations of the input function values $\mathbf{f}_{i}(\mathbf{x})\in\mathbb{F}({\Omega};^{p_{f}})$ and the corresponding output function values $\mathbf{u}_{i}(\mathbf{x})\in\mathbb{U}({\Omega};^{p_{u}})$. Here, $i$ denotes the sample index, ${\Omega}\in^{p}$ is the bounded domain of interest, and $\mathbb{F}$ and $\mathbb{U}$ describe the Banach spaces of functions taking values in ${}^{p_{f}}$ and ${}^{p_{u}}$, respectively. To model the physical laws of such a system, we aim to learn the intrinsic operator $\mathcal{G}:\mathbb{F}\rightarrow\mathbb{U}$, that maps the input function $\mathbf{f}(\mathbf{x})$ to the output function $\mathbf{u}(\mathbf{x})$. In other words, given the measurements of $\mathbf{f}_{i}$ and $\mathbf{u}_{i}$ on a collection of points $\chi=\{\mathbf{x}_{j}\}^{M}_{j=1}\subset{\Omega}$, we seek to learn the physical response by constructing a parameterized surrogate operator of $\mathcal{G}$: $\tilde{\mathcal{G}}[\mathbf{f};\theta](\mathbf{x})\approx\mathbf{u}(\mathbf{x})$, where the trainable parameter set $\theta$ is obtained by solving the optimization problem:
$\min_{\theta\in\Theta}\mathcal{L}(\theta):=$ $\min_{\theta\in\Theta}\sum_{i=1}^{N}\dfrac{1}{N}[C(\tilde{\mathcal{G}}[\mathbf{f}_{i};\theta],\mathbf{u}_{i})]$ (1)
$\approx$ $\min_{\theta\in\Theta}\dfrac{1}{N}\sum_{i=1}^{N}\dfrac{\sum_{j=1}^{M}{\left|\left|\tilde{\mathcal{G}}[\mathbf{f}_{i};\theta](\mathbf{x}_{j})-\mathbf{u}_{i}(\mathbf{x}_{j})\right|\right|}^{2}}{\sum_{j=1}^{M}{\left|\left|\mathbf{u}_{i}(\mathbf{x}_{j})\right|\right|}^{2}}\text{ .}$
@@
Although our approach is agnostic to any NO architecture, for the purpose of demonstration we consider the generic integral neural operators (Li et al., 2020b, c; You et al., 2022a, b) as the base models. In this context, an $L$-layer NO has the following form:
$\tilde{\mathcal{G}}[\mathbf{f};\theta](\mathbf{x}):=\mathcal{Q}\circ\mathcal{J}_{L}\circ\cdots\circ\mathcal{J}_{1}\circ\mathcal{P}[\mathbf{f}](\mathbf{x})\text{ ,}$ (2)
@@
$\mathbf{h}$ $(\mathbf{x},l+1)=\mathcal{J}^{FNO}_{l}[\mathbf{h}(\mathbf{x},l)]$ (3)
::=:= $\sigma(W_{l}\mathbf{h}(\mathbf{x},l)+\mathcal{F}^{-1}[A_{l}\cdot\mathcal{F}[\mathbf{h}(\cdot,l)]](\mathbf{x})+\mathbf{c}_{l})\text{ ,}$
@@ 3.1 Built-In Divergence-Free Prediction
$\dfrac{\partial\rho}{\partial t}(\mathbf{x},t)+\nabla_{\mathbf{x}}\cdot\mu(\mathbf{x},t)=0\text{ ,}\quad\text{ for }(\mathbf{x},t)\in{\Omega}\times[0,T]\text{ ,}$ (5)
@@
where $\rho:^{p+1}\rightarrow^{+}$ is the volume density of the quantity to be conserved, $\mu:^{p+1}\rightarrow^{p}$ is the flux describing how this quantity flows, and $\nabla_{\mathbf{x}}$ denotes the gradient operator with respect to the spatial dimensions. Equation 5 states that the amount of the conserved quantity within a volume can only change by the amount of the quantity that flows in and out of the volume. Taking the mass conservation law as an instance, $\rho$ is the fluid’s density, and $\mathbf{u}$ is the velocity at each point, and then the mass flux becomes $\mu=\rho\mathbf{u}$. When considering the (quasi-)static problem, i.e., $\rho$ is independent of time, we have $\nabla_{\mathbf{x}}\cdot\mathbf{u}=0$, where the divergence is taken only over spatial variables. Such a divergence-free equation is also referred to as the condition of incompressibility in fluid dynamics. Moreover, we note that (5) can always be expressed as a divergence-free equation when considering both spatial and temporal dimensions. As a matter of fact, by taking $\mathbf{X}=(\mathbf{x},t)$ and $\mathbf{U}=(\mu,\rho)$, (5) can be re-written as $\nabla_{\mathbf{X}}\cdot\mathbf{U}=0$, there the gradient operator $\nabla_{\mathbf{X}}$ is taken with respect to both $\mathbf{x}$ and $t$. Then, the conservation law becomes equivalent to the static case, if we consider the augmented variable $\mathbf{X}$ on the augmented domain $\tilde{{\Omega}}:={\Omega}\times[0,T]$. Hence, in the following derivation we focus on the static case, with the goal of designing an NO architecture with the continuity equation $\text{div}(\mathbf{u}):=\nabla_{\mathbf{x}}(\mathbf{u})=0$ automatically guaranteed for its output function $\mathbf{u}$.
@@
To construct divergence-free NOs, we start from differential forms in p, following the derivations in e.g., Barbarosie (2011); Kelliher (2021). More details are provided in Appendix A. Generally, an object that may be integrated over a $k-$dimensional manifold is called a $k-$form, which can be expressed as $\mu=\sum_{1\leq i_{1},i_{2},\cdots,i_{k}\leq p}\mu_{(i_{1},\cdots,i_{k})}dx_{i_{1}}\wedge\cdots\wedge dx_{i_{k}}:=\sum_{I}\mu_{I}dx_{I}$. For instance, any scalar function $g(\mathbf{x})\in C^{\infty}$ is a $0-$form, and $v=\sum_{i=1}^{p}v_{i}(\mathbf{x})dx_{i}$, $v_{i}(\mathbf{x})\in C^{\infty}$, is a $1-$form. Denote $d$ as the exterior derivative, which is an operation acting on a $k-$form and produces a $(k+1)-$form (e.g., $dg(\mathbf{x})=\sum_{i=1}^{p}\frac{\partial g}{\partial x_{i}}dx_{i}$), and the Hodge operator as $\star$, which matches to each $k-$form an $(p-k)-$form via $\star dx_{I}:=(-1)^{\sigma(I,I^{c})}dx_{I^{c}}$. Here, $I^{c}:=\{1,\cdots,n\}\backslash I$ is the complement set of $I$, and $\sigma(I,I^{c})$ is the sign of the permutation $(I,I^{c})$. A simple calculation yields that $\star\star\mu=(-1)^{k(p-k)}\mu$ and $dd\mu=0$ for any $k-$form $\mu$.
@@
To see the connection of these definitions to our goal, we note that our output function, $\mathbf{u}:^{p}\rightarrow^{p}$, can be equivalently expressed as a $1-$form $\mathbf{u}=\sum^{p}_{i=1}u_{i}(\mathbf{x})dx_{i}$. Then, the divergence of $\mathbf{u}$ can be expressed as $d\star\mathbf{u}$, and the above fundamental properties of the exterior derivative yields $\text{div}(\star d\mu)=d\star\star d\mu=dd\mu=0$ for any $(p-2)-$form $\mu$. That means, any vector field $\mathbf{u}$ satisfying $\mathbf{u}=\star d\mu$ is divergence-free. Our goal is therefore to parameterize the output function of NOs in the form of $\star d\mu$, or equivalently, as $\star d(\sum_{1\leq i_{1},\cdots,i_{p-2}\leq p}\mu_{(i_{1},\cdots,i_{p-2})}dx_{i_{1}}\wedge\cdots\wedge dx_{i_{p-2}}))=\sum_{i=1}^{p}\sum_{j=1}^{p}\frac{\partial\mu_{ij}}{\partial x_{j}}dx_{i}$. Here, $\mu_{ij}(\mathbf{x})$ is a function defined on ${\Omega}$ with $\mu_{ij}(\mathbf{x})=\mu_{(i,j)^{c}}$ for $i<j$ and $\mu_{ij}(\mathbf{x})=-\mu_{(j,i)^{c}}$ otherwise, as real-valued alternating functions. That means, taking any $\mu:=[\mu_{ij}(\mathbf{x})]$ satisfies $\mu_{ij}=-\mu_{ji}$ (a $p$ by $p$ skew-symmetric matrix-valued function), we have $\text{div}(\star d\mu)=0$, where $\star d\mu=[\text{div}(\mu_{1}),\cdots,\text{div}(\mu_{p})]^{T}$ and $\mu_{i}$ stands for the $i-$th row of $\mu$.